OpenStudy (refrac532):

Rewrite the function y = (3.1)(2/3)^-x in the form y = ab^x.

ILovePuppiesLol (ilovepuppieslol):

@imqwerty

OpenStudy (refrac532):

That doesn't really help @ILovePuppiesLol

OpenStudy (mathmale):

Task: write y = (3.1)(2/3)^-x in the form y = ab^x 1. Focus on the exponential function part: \[(\frac{ 2 }{ 3 })^{-x}\]

OpenStudy (mathmale):

Please use either Equation Editor or the Draw utility to present the exponentiation properly: See my expression immediately above, or 

OpenStudy (mathmale):

You are asked to write this with a positive exponent instead of a negative one. There are several ways in which you could accomplish that. One would be to simply invert (turn upside down) the (2/3) and then use +x as exponent instead of -x. Try this, please. Share your result.

OpenStudy (refrac532):

Would the answer be\[y=(3.1)(\frac{ 3 }{ 2 })^x\]

OpenStudy (refrac532):

I don't know, I feel like that's wrong, but I think it's right...

OpenStudy (refrac532):

@mathmale

OpenStudy (anonymous):

correct.

OpenStudy (refrac532):

Alright, thanks!